(D2+7DD'+12D'2)z=sinhx

1
Expert's answer
2021-05-07T10:31:26-0400

"(D^2+7DD'+12D'^2)z=sinhx"


Auxilary equation is-

  "m^2+7m+12=0\\

 \Rightarrow(m+3)(m+4)=0\\

 \Rightarrow m=-3,-4"


Complimentary function CF is-

       "CF= f_1(y-3x)+f_2(y-4x)"


Then Paticular Integral-


         "PI=\dfrac{1}{D^2+7DD'+12D'^2}sinhx"

     

           "=\dfrac{1}{D^2+7DD'+12D'^2}(\dfrac{e^x-e^{-x}}{2})"


           "=\dfrac{1}{2}(\dfrac{e^x}{D^2+7DD'+12D'^2}-\dfrac{e^{-x}}{D^2+7DD'+12D'^2})"


          "=\dfrac{1}{2}(\dfrac{e^x}{(1^2+7(1)+12(0)}-\dfrac{e^{-x}}{(-1)^2+7(1)+12(0)})"


         "=\dfrac{1}{2}(\dfrac{e^x-e^{-x}}{8})"


         "=\dfrac{sinhx}{8}"


  Solution is-

 "z=f_1(y+3x)+f_2(y+6x)+\dfrac{sinhx}{8}"